what is the value of sinA+cos²A-1 if 5cosecA=7.Please answer this question as quick as you can.

Grade:10

1 Answers

Arun
5 years ago
As 5 cosecA = 7
.hence
sin A = 5/7
Now cos²A = 1 - sin²A = 1 - 25/49 = 24/49
Hence
SinA + cos² A -1 = 5/7 + 24/49 -1 = (35 +24 - 49)/49 = 10/49